OpenStudy (anonymous):

y-5=-8...solve and show all steps....my son has -14-5=-8....y = -14

OpenStudy (anonymous):

-13 sorry

OpenStudy (chihiroasleaf):

y-5 = -8 y = -8 + 5 y = -3

OpenStudy (anonymous):

y=-3 you add +5 to both sides to get y alone

OpenStudy (anonymous):

wow...he missed that one....thank you...i'll explain it to him

OpenStudy (chihiroasleaf):

you're welcome..